Radius=10cm Time=45sec Uniform speed=????????????????????????????????????????????????????

Radius=10cm
Time=45sec
Uniform speed=????????????????????????????????????????????????????

Grade:9

3 Answers

Arun
3 years ago
Dear student
 
Question is not clear. Please post proper question or attach an image. I will be happy to help you
Jyothi sri
3 years ago
radius is 10cm = 0.1 m
so, circumference = 2*22/7*0.1= 44/70 m
now, the distance covered by the object is the circumference as it is not mentioned about number of rounds. so it is assumed that the object moved for 1 round. using speed= distance/time formula , we get
speed= 44/70 *1/45= 0.0139 m/s {approx.}
Ayush_Deep
3 years ago
Since the particle moves with constant speed.
Total time taken to complete one revolution is 45 sec
Total distance traveled in one revolution is=2πr
Uniform speed = 2πr/45
=      20π/45 cm/sec
=     1.39 cm/sec
Hope you got your answer